OpenStudy (anonymous):

the areas of two equilateral triangles are 27 yd squared and 75 yd squared. Find the ratio of their perimeters.

OpenStudy (anonymous):

first you would have to find the Perimeters of each triangle.

OpenStudy (anonymous):

since you have equilateeral triangles all of the three sides would be equal. You can extract the sides from the formula since you know the Area.

OpenStudy (anonymous):

ok thank you I got it from here (:

OpenStudy (anonymous):

np gl

OpenStudy (phi):

they want you to know this useful fact if the ratio of the sides of two similar shapes is \( \frac{s_1}{s_2}\), the ratio of their areas will be \( \frac{s_1^2}{s_2^2}\) In other words, the areas will be the ratio of sides squared for an equilateral triangle, the 3 sides are equal and the perimeter is 3 s we can say \[ \frac{9 s_1^2}{9 s_2^2} = \frac{27}{75} \] if we take the square root of both sides we will find the ratio of their perimeters \[ \frac{\sqrt{27}}{\sqrt{75}}= \frac{3 \sqrt{3}}{5\sqrt{3}}= \frac{3}{5}\]